Calculate the work done by an applied force of 15 N [35° S of W] on an object that has moved 7.5 m [N 37° W].
1
Expert's answer
2020-05-15T08:42:45-0400
"W=Fs\cos{\alpha}"

"W=(15)(7.5)\cos{(37+35)}=34.8\ J"